Consider the combustion of the wax C_27H_56 (380 g
Answer #1

C27H56 + O2 --> CO2 + H2O

Balance it

C27H56 + 41O2 --> 27CO2 + 28H2O

mol = mass/MW = 500/380 = 1.315 mol of wax

1 mol of wax = 27 mol of CO2

then

1.315 = 1.315*27 = 35.505 mol of CO2

in STP

1 mol = 22.4 L

then 35.505*22.4 = 795.312 L of CO2

answer is C